OpenStudy (anonymous):

Graph the equation with a diameter that has endpoints at (-3, 4) and (5, -2). Label the center and at least four points on the circle. Write the equation of the circle. @dan815 could you please help

OpenStudy (anonymous):

i just need the equation

OpenStudy (anonymous):

hey @highjumpcollect I'm sorry I can't help. I can graph this it's very easy but I wouldn't know the equation...

OpenStudy (anonymous):

ok well i just need the equation

OpenStudy (anonymous):

all right then sorry I wish I could help I'm sure someone else will come soon :)

OpenStudy (anonymous):

@Mindblast3r @Mimi_x3 @dan815

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

When you put the center of the circle on the graph, what did you give as its coordinates?

Directrix (directrix):

The answer for the circle equation will be in the format of the standard form shown in the attachment.

Directrix (directrix):

If you don't participate, I cannot help you.

OpenStudy (anonymous):

yes sorry i was working on it.. i can't figure the radius out.. @Directrix

OpenStudy (anonymous):

I just need to find the radius @Directrix

Directrix (directrix):

Let's try this. Get the midpoint of the segment with these endpoints (-3, 4) and (5, -2)

Directrix (directrix):

Get the average of the x terms and the average of the two y-coordinates. Post what you get.

OpenStudy (anonymous):

(1,1)

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

please help me @Directrix

Directrix (directrix):

Okay with the center. We need one more thing, the radius. You can get it by finding the distance between (1,1) and (-3, 4). Then, we will fit that into the equation and be finished.

OpenStudy (anonymous):

how is that done?

OpenStudy (anonymous):

I got 5 is that correct?

Directrix (directrix):

Use the distance formula

Directrix (directrix):

Yes, 5 is correct.

OpenStudy (anonymous):

YAY thank you SO much!!!!

Directrix (directrix):

Now, use this to get the equation.

Directrix (directrix):

So, what is the equation?

OpenStudy (anonymous):

Yes, thank u

Directrix (directrix):

No, we have to get the equation.

Directrix (directrix):

(x - 1)^2 + what = what

OpenStudy (anonymous):

yes i already got it (x-1)^2+y-1)^2= 25

Directrix (directrix):

(x-1)^2 + (y-1)^2 = 5^2